OpenStudy (kweenlexis):

Which statements are true about the ordered pair (−1,−4) and the system of equations? x−y=3 7x−y=−3 Select each correct answer. A. The ordered pair ​ (−1,−4) ​ is a solution to the system of linear equations. B. The ordered pair ​(−1,−4) ​ is not a solution to the system of linear equations. C. When ​ (−1,−4) ​ is substituted into the second equation, the equation is true. D. When ​ (−1,−4) ​ is substituted into the first equation, the equation is false. E. When ​ (−1,−4) ​ is substituted into the second equation, the equation is false

OpenStudy (kweenlexis):

@mathstudent55

OpenStudy (mathstudent55):

Oh, I see, you corrected the equations. All you need to do is replace x with -1 and y with -4 in each equation and see if it makes true statements.

OpenStudy (mathstudent55):

The first equation is: x - y = 3 Now we substitute x with -1 and y with -4: -1 - (-4) = 3 Can you do the subtraction on the left side?

OpenStudy (mathmale):

You have two basic choices here: 1) Actually solve the system yourself. 2) Pick points, one at a time, from the given answer choices, and subst. the given x- and y-values into BOTH of the given equations. Are both equations then true? Then you've found a solution. Not true? Then that point does not represent a solution. Personally, I'd prfer to solve the system of equations myself, to be sure which answer is the correct solution to this system.

OpenStudy (kweenlexis):

It's true?

OpenStudy (mathstudent55):

What is -1 - (-4) = ?

OpenStudy (kweenlexis):

3

OpenStudy (kweenlexis):

and 7*-1-(-4)=-3

OpenStudy (kweenlexis):

so its true

OpenStudy (kweenlexis):

so A,C and E?

OpenStudy (kweenlexis):

@mathstudent55